# Subdiffusive concentration for the chemical distance in Bernoulli percolation

### v1

@@ 3. Lattice animals of dependent weight
In the previous section, we derive that the impact of resampling an edge along a modified geodesic can be controlled by its effective radius. Our current goal is to estimate the total cost of rampling, i.e., the sum of the effective radii along a random set. Fortunately, although these effective radii are not independent, they exhibit relatively local dependence (Remark 2.3) and furthermore, satisfy the exponentially decaying tail probability (Proposition 2.4). Moreover, the theory of greedy lattice animals can help us control the total weight of paths in locally dependent environments.
@@
    $(I_{e,M})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ are $aM-$dependent, i.e. for all $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$, the variable $I_{e,M}$ is independent of all variables $(I_{e^{\prime},M})_{e^{\prime}\not\in\Lambda_{aM}(e)}$.
@@
$\Xi_{L}:=\{\gamma:\gamma\text{ is a set of edges in }\Lambda_{L};\,|\gamma|\leq L\}.$
@@
[CN19, Lemma 2.6] Let $M\geq 1,a,A>0$ and $(I_{e,M})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ be a collection of random variables satisfying (E1) and (E2). Then, there exists a positive constant $C=C(a,A,d)$ such that
@@
    If $t\geq CM^{d}\max\left(1,MLq_{M}^{1/d}\right)$, then
@@
In [CN19, Lemma 2.6], although the result is stated for site-percolation, it also holds true for edge-percolation by similar arguments, see for example [CNN24, Lemma 2.6].
@@
    Let $\gamma$ be a random path starting from 00 in the same probability space of $X$. Then for all $L\geq 1$,
@@
    Let $m\geq 1$ and $\gamma$ be a random set of edges of $\Lambda_{m}$ in the same probability space of $X$. Then for all $L\geq 1$,
@@
    Let $L\geq 1$ and $K\in{\mathbb{N}}$ such that $LM\phi(M)^{1/d}\geq 1$ for all $1\leq M\leq K$. Assume further that $X_{e}\leq K$ for all $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$. There exist constants $C_{1},c_{1}>0$ such that for any random path $\gamma$ starting from 00 in the same probability space of $X$,
@@
${\mathbb{E}}\left[\left(\max_{\gamma\in\Xi_{L}}\sum_{e\in\gamma}f(X_{e})\right)^{2}\right]={\mathbb{E}}\left[\max_{\gamma\in\Xi_{L}}\left(\sum_{e\in\gamma}f(X_{e})\right)^{2}\right]\leq{\mathbb{E}}\left[\max_{\gamma\in\Xi_{L}}|\gamma|\sum_{e\in\gamma}f^{2}(X_{e})\right]\leq L{\mathbb{E}}\left[\max_{\gamma\in\Xi_{L}}\sum_{e\in\gamma}f^{2}(X_{e})\right],$ (3.1)
@@
$A^{\gamma}_{M}:=\left|\{e\in\gamma:M-1<X_{e}\leq M\}\right|=\sum_{e\in\gamma}I_{e,M},$
@@
$\sum_{e\in\gamma}f^{2}(X_{e})\leq\sum_{M\geq 1}f^{2}_{*}(M)A^{\gamma}_{M}+f^{2}(0)|\gamma|,$ (3.2)
@@
$=\sum_{M\geq 1}f_{*}^{2}(M){\mathbb{E}}\left[N_{L,M}\right]+f^{2}(0)L,$ (3.3)
@@
using (H). Finally, combining the above estimate with (3.1), we obtain (i). For (ii), we decompose
@@
$\leq{\mathcal{O}}(L^{2})+\sum_{\ell=L}^{\infty}\ell{\mathbb{E}}\left[\max_{\gamma\in\Xi_{\ell}}\sum_{e\in\gamma}f^{2}(X_{e}){\mathbb{I}}(|\gamma|=\ell)\right],$ (3.5)
by using (i). Moreover, by Cauchy-Schwarz inequality and (i),
${\mathbb{E}}\left[\max_{\gamma\in\Xi_{\ell}}\sum_{e\in\gamma}f^{2}(X_{e}){\mathbb{I}}(|\gamma|=\ell)\right]$ $\leq{\mathbb{E}}\left[\left(\max_{\gamma\in\Xi_{\ell}}\sum_{e\in\gamma}f^{2}(X_{e})\right)^{2}\right]^{1/2}{\mathbb{E}}\left[{\mathbb{I}}(|\gamma|=\ell)\right]^{1/2}$
@@
Finally, we show (iv). Using $X_{e}\leq K$ and a similar estimate as in (3.2), we have for any $C_{1}\geq 2f(0)$,
@@
$\leq\mathbb{P}\left(\sum_{M=1}^{K}f_{*}(M)N_{L,M}\geq C_{1}L/2\right)+\mathbb{P}\left(|\gamma|\geq L\right).$ (3.6)
@@
Finally, we shall apply Lemma 3.3 to the sequence of truncated effective radii defined in (2.4): $\hat{R}_{e}=\min\{C_{*}R_{e},\log^{2}n\}$ for $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$.
@@
    Let $\gamma$ be a random path in the same probability space of $(R_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$, starting from 00 and satisfying ${\mathbb{P}}(|\gamma|=\ell)\leq C\ell^{-7}$ for all integer $\ell\geq Cn$. Then
    ${\mathbb{E}}\Big{[}\Big{(}\sum_{e\in\gamma}\hat{R}_{e}\Big{)}^{2}\Big{]}\leq{\mathbb{E}}\Big{[}\Big{(}\sum_{e\in\gamma}\hat{R}^{2}_{e}\Big{)}^{2}\Big{]}\leq Cn^{2}.$
@@
    Let $\gamma$ be a random set of edges of $\Lambda_{m}$ for some integer $m\geq 1$ in the same probability space of $(R_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ satisfying ${\mathbb{P}}(|\gamma|=\ell)\leq C\ell^{-7}$ for all integer $\ell\geq Cm$. Then
    ${\mathbb{E}}\Big{[}\sum_{e\in\gamma}\hat{R}_{e}^{2}\Big{]}\leq Cm.$
@@
    For any $c>0$, there exists $C_{1}=C_{1}(c)$ such that for all $\gamma$, a random path starting from 00 in the same probability space of $(R_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$,
@@
Let $f(x)=x^{2}$ and $\phi(x)=x^{-d^{2}+11d}$, so (H) is verified. By Remark 2.3 and Proposition 2.4, the radii $(\hat{R}_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ satisfy (P1) and (P2). Thus, the parts (i) and (ii) follow from Lemma 3.3 (ii) and (iii), respectively. Finally, applying Lemma 3.3 (iv) to $L=cn,K=\lceil\log^{2}n\rceil$, we obtain (iii). ∎
@@ 4. Subdiffusive concentration of ${\mathrm{T}}_{n}$
The proof strategy for Theorem 1.3 relies on establishing a connection between bounds on $\operatorname{Var}(e^{\lambda{\mathrm{T}}_{n}/2})$ and exponential concentration (see [BR08, Lemma 4.1]). To achieve the required variance bound, we employ the Falik-Samorodnitsky inequality (Lemma 4.1) to a martingale decomposition of the random variable $e^{\lambda{\mathrm{F}}_{m}}$, where ${\mathrm{F}}_{m}$ represents an averaged version of the passage time. This approach was first introduced by Benaïm and Rossignol in [BR08] and subsequently utilized by Damron, Hanson, and Sosoe in [DHS14]. Finally, we estimate the tails of the true passage time ${\mathrm{T}}_{n}$ based on those of ${\mathrm{F}}_{m}$.
@@ 4.1. Variance bound via entropy inequality
$\forall i\geq 1,\quad\Delta_{i}:={\mathbb{E}}[G\mid{\mathcal{F}}_{i}]-{\mathbb{E}}[G\mid{\mathcal{F}}_{i-1}]={\mathbb{E}}[G(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-G(t_{e_{i}},t_{e_{i}^{c}})\mid{\mathcal{F}}_{i-1}],$
where $t^{\prime}_{e_{i}}$ is an independent copy of $t_{e_{i}}$. We will bound the variance of $G$ based on the following entropy inequality by Falik and Samorodnitsky [FS07, Lemma 2.3].
@@
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]\geq\operatorname{Var}[G]\log\dfrac{\operatorname{Var}[G]}{\sum_{i=1}^{\infty}({\mathbb{E}}[|\Delta_{i}|])^{2}},$ (4.1)
@@
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]\leq C\sum_{i=1}^{\infty}{\mathbb{E}}[(G(b,t_{e_{i}^{c}})-G(a,t_{e_{i}^{c}}))^{2}].$ (4.2)
4.2. Proof of Theorem 1.3
Instead of directly showing the subdiffusive concentration of ${{\mathrm{T}}_{n}}$, we will employ a strategy inspired by Benjamini, Kalai, and Schramm in [BKS03], known as the BKS trick. This approach involves proving the subdiffusive concentration for a geometric average of passage time, a concept introduced in [BKS03] and later used in [AZ13, DHS14, DHS15, BDG20]. It is expected that the majority of edges in the lattice have a low probability of lying in the geodesic of ${\mathrm{T}}_{n}$, meaning they have a small influence. However, this does not hold for edges very close to the origin, and the BKS trick provides a way to circumvent this issue.
First of all, let us define a spatial average of the first passage time,
${{\mathrm{F}}_{m}}:=\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}{\mathrm{T}}_{z},$ (4.3)
@@
$\forall z\in{\mathbb{Z}}^{d},\quad{\mathrm{T}}_{z}:={\mathrm{T}}(z,z+n{\mathbf{e}}_{1}),\qquad m:=\lfloor n^{1/4}\rfloor.$
@@
$\forall\,|\lambda|<\dfrac{1}{\sqrt{K}},\quad\operatorname{Var}\left[e^{\lambda{{\mathrm{F}}_{m}}}\right]\leq K\lambda^{2}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]<\infty,$ (4.4)
@@
The following result is a direct consequence of Theorem 4.3. We refer the reader to [BR08, Lemma 4.1] for a proof.
@@
$\mathbb{P}\left(\left|{{\mathrm{F}}_{m}}-{\mathbb{E}}[{{\mathrm{F}}_{m}}]\right|\geq\sqrt{\frac{n}{\log n}}\kappa\right)\leq c^{-1}\exp(-c\kappa)\quad\forall\kappa\geq 0.$ (4.5)
@@
where we recall that $z^{*}$ is the closest point of $z$ in the infinite cluster ${\mathcal{C}}_{\infty}$. The last two terms are bounded by $c_{1}^{-1}\exp\left(\tfrac{-c_{1}t}{\log^{2}n}\right)$, for some positive constants $c_{1}$ using Lemma 2.8, whereas by Lemma 2.7, the first term is bounded by $c_{2}^{-1}\exp(-c_{2}t)$ when $t\geq C\|x-y\|_{\infty}$ with some $C,c_{2}>0$. Hence, the result follows. ∎
@@
$\leq|{{\mathrm{F}}_{m}}-{\mathbb{E}}[{{\mathrm{F}}_{m}}]|+|{{\mathrm{T}}_{n}}-{{\mathrm{F}}_{m}}|.$ (4.7)
@@
$\mathbb{P}\left(|{{\mathrm{T}}_{n}}-{\mathbb{E}}[{{\mathrm{T}}_{n}}]|\geq 4M\right)\leq\mathbb{P}\left(|{{\mathrm{F}}_{m}}-{\mathbb{E}}[{{\mathrm{F}}_{m}}]|\geq 2M\right)+\mathbb{P}\left(|{{\mathrm{T}}_{n}}-{{\mathrm{F}}_{m}}|\geq 2M\right).$ (4.8)
@@
$|{{\mathrm{T}}_{n}}-{{\mathrm{F}}_{m}}|=\Big{|}{{\mathrm{T}}_{n}}-\dfrac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}{\mathrm{T}}_{z}\Big{|}$ $\leq\dfrac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}|{\mathrm{T}}(0,n{\mathbf{e}}_{1})-{\mathrm{T}}(z,z+n{\mathbf{e}}_{1})|$
$\leq\dfrac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}({\mathrm{T}}(0,z)+{\mathrm{T}}(n{\mathbf{e}}_{1},n{\mathbf{e}}_{1}+z)).$ (4.9)
@@
$\max_{z\in\Lambda_{m}}{\mathrm{T}}(0,z)\geq M\text{ or }\max_{z\in\Lambda_{m}}{\mathrm{T}}(n{\mathbf{e}}_{1},n{\mathbf{e}}_{1}+z)\geq M.$ (4.10)
@@
$\qquad\qquad\qquad=2\mathbb{P}\left(\max_{z\in\Lambda_{m}}{\mathrm{T}}(0,z)\geq M\right)\leq 2|\Lambda_{m}|\max_{z\in\Lambda_{m}}\mathbb{P}\left({\mathrm{T}}(0,z)\geq M\right),$ (4.11)
@@
$\max_{z\in\Lambda_{m}}\mathbb{P}\left({\mathrm{T}}(0,z)\geq M\right)\leq c^{-1}\exp(-cM/\log^{2}n),$ (4.12)
@@
Combining this with Corollary 4.4 and (4.8), it follows that
$\mathbb{P}\left(|{{\mathrm{T}}_{n}}-\mathbb{E}[{{\mathrm{T}}_{n}}]|\geq\sqrt{\frac{n}{\log n}}\kappa\right)\leq c^{-1}\exp(-c\kappa)\qquad\forall\kappa\geq 0,$ (4.13)
@@
Proof of Theorem 4.3. According to the Lemma 4.1, the variance bound is induced by two crucial factors: the estimate of total influence and total entropy. These keys are presented in the following results.
@@
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq C\lambda^{2}n{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\quad\forall\,|\lambda|\leq\tfrac{1}{\log^{2(d+11)}n}.$ (4.14)
@@
$\operatorname{Var}\big{[}e^{\lambda{\mathrm{F}}_{m}}\big{]}\leq C\left(\log\frac{\operatorname{Var}[e^{\lambda{\mathrm{F}}_{m}}]}{C\lambda^{2}n^{(9-d)/8}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]}\right)^{-1}\lambda^{2}n{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}.$ (4.15)
We can assume that
$\operatorname{Var}[e^{\lambda{\mathrm{F}}_{m}}]\geq C\lambda^{2}n^{15/16}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right],$ (4.16)
since otherwise there is nothing left to prove. By (4.15) and (4.16), there exists a constant $c>0$ such that for any $|\lambda|\leq\tfrac{1}{\log^{2(d+11)}n}$,

### v2

@@ 3. Lattice animals of dependent weight
In the previous section, we derive that the impact of resampling an edge along a modified geodesic can be controlled by its effective radius. Our current goal is to estimate the total cost of resampling, i.e., the sum of the effective radii along a random set. Fortunately, although these effective radii are not independent, they exhibit relatively local dependence (Remark 2.5) and furthermore, satisfy the exponentially decaying tail probability (Proposition 2.6). Moreover, the theory of greedy lattice animals can help us control the total weight of paths in locally dependent environments.
@@
    $(I_{e,M})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ are aMaM-dependent, i.e. for all $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$, the variable $I_{e,M}$ is independent of all variables $(I_{e^{\prime},M})_{e^{\prime}\not\in\Lambda_{aM}(e)}$.
@@
$\Xi_{L}:=\{\gamma:\gamma\text{ is a path in }\Lambda_{L};\,|\gamma|\leq L\}.$
@@
[8, Lemma 2.6] Let $M\geq 1,a,A>0$ and $(I_{e,M})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ be a collection of random variables satisfying (E1) and (E2). Then, there exists a positive constant $C=C(a,A,d)$ such that
@@
    If $t\geq CM^{d}\max\left\{1,MLq_{M}^{1/d}\right\}$, then
@@
In [8, Lemma 2.6], although the result is stated for site-percolation, it also holds true for edge-percolation by similar arguments, see for example [9, Lemma 2.6].
@@
    Let $\gamma$ be a random path starting from 0 in the same probability space as $X$. Then for all $L\geq 1$,
@@
    Let $m\geq 1$ and $\gamma$ be a random path with ${\bf s}(\gamma)\in\Lambda_{m}$ in the same probability space as $X$. Then for all $L\geq 1$,
@@
    Let $L\geq 1$ and $K\in{\mathbb{N}}$ such that $LM\phi(M)^{1/d}\geq 1$ for all $1\leq M\leq K$. Assume that $X_{e}\leq K$ for all $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$. There exist constants $C_{1},c_{1}>0$ such that for any random path $\gamma$ starting from 0 in the same probability space as $X$,
@@
${\mathbb{E}}\left[\left(\max_{\gamma\in\Xi_{L}}\sum_{e\in\gamma}f(X_{e})\right)^{2}\right]={\mathbb{E}}\left[\max_{\gamma\in\Xi_{L}}\left(\sum_{e\in\gamma}f(X_{e})\right)^{2}\right]\leq{\mathbb{E}}\left[\max_{\gamma\in\Xi_{L}}|\gamma|\sum_{e\in\gamma}f^{2}(X_{e})\right]\leq L{\mathbb{E}}\left[\max_{\gamma\in\Xi_{L}}\sum_{e\in\gamma}f^{2}(X_{e})\right],$ (36)
@@
$A^{\gamma}_{M}:=\left|\{e\in\gamma:M-1\leq X_{e}<M\}\right|=\sum_{e\in\gamma}I_{e,M},$
@@
$\sum_{e\in\gamma}f^{2}(X_{e})\leq\sum_{M\geq 1}f^{2}_{*}(M)A^{\gamma}_{M}+f^{2}(0)|\gamma|,$ (37)
@@
$=\sum_{M\geq 1}f_{*}^{2}(M){\mathbb{E}}\left[N_{L,M}\right]+f^{2}(0)L,$ (38)
@@
where the second equality follows from
$\sum_{M\geq 1}f_{*}^{2}(M)M^{d+1}\phi(M)^{1/d}<\infty,$
which is guaranteed by (H). Finally, combining the above estimate with (36), we obtain (i).
For (ii), we decompose
@@
$\leq{\mathcal{O}}(L^{2})+\sum_{\ell=L}^{\infty}\ell{\mathbb{E}}\left[\max_{\sigma\in\Xi_{\ell}}\sum_{e\in\sigma}f^{2}(X_{e}){\mathbb{I}}(|\gamma|=\ell)\right],$ (40)
by using (i). Moreover, by first using Cauchy-Schwarz inequality and then applying (i) to $f^{2}$, we obtain
${\mathbb{E}}\left[\max_{\sigma\in\Xi_{\ell}}\sum_{e\in\sigma}f^{2}(X_{e}){\mathbb{I}}(|\gamma|=\ell)\right]$ $\leq{\mathbb{E}}\left[\left(\max_{\sigma\in\Xi_{\ell}}\sum_{e\in\sigma}f^{2}(X_{e})\right)^{2}\right]^{1/2}{\mathbb{E}}\left[{\mathbb{I}}(|\gamma|=\ell)\right]^{1/2}$
@@
Finally, we show (iv). Using $X_{e}\leq K$ and a similar estimate as in (37), we have for any $C_{1}\geq 2f(0)$,
@@
$\leq\mathbb{P}\left(\sum_{M=1}^{K}f_{*}(M)N_{L,M}\geq C_{1}L/2\right)+\mathbb{P}\left(|\gamma|\geq L\right).$ (41)
@@
Finally, we shall apply Lemma 3.3 to the sequence of truncated effective radii defined in (20): $\hat{R}_{e}=\min\{C_{*}R_{e},\log^{2}n\}$ for $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$.
@@
    Let $\gamma$ be a random path in the same probability space as $(R_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$, starting from 0 and satisfying ${\mathbb{P}}(|\gamma|=\ell)\leq C\ell^{-7}$ for all integers $\ell\geq Cn$. Then
    ${\mathbb{E}}\left[\left(\sum_{e\in\gamma}\hat{R}_{e}\right)^{2}\right]\leq{\mathbb{E}}\left[\left(\sum_{e\in\gamma}\hat{R}^{2}_{e}\right)^{2}\right]\leq Cn^{2}.$
@@
    Let $\gamma$ be a random path in the same probability space as $(R_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ with ${\bf s}(\gamma)\in\Lambda_{m}$ for some integer $m\geq 1$. Suppose that ${\mathbb{P}}(|\gamma|=\ell)\leq C\ell^{-7}$ for all integers $\ell\geq Cm$. Then
    ${\mathbb{E}}\left[\sum_{e\in\gamma}\hat{R}_{e}^{2}\right]\leq Cm.$
@@
    For any $c>0$, there exists $C_{1}=C_{1}(c)$ such that for all $\gamma$, a random path starting from 0 in the same probability space as $(R_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$,
@@
Let $f(x)=x^{2}$ and $\phi(x)=x^{-d^{2}-11d}$, so (H) is verified. By Remark 2.5 and Proposition 2.6, the radii $(\hat{R}_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ satisfy (P1) and (P2). We take $X_{e}=\hat{R}_{e}$ for each $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$. Thus, the parts (i) and (ii) follow from Lemma 3.3 (ii) and (iii), respectively. Finally, applying Lemma 3.3 (iv) to $L=cn,K=\lceil\log^{2}n\rceil$, we obtain (iii). ∎
@@ 4. Subdiffusive concentration of ${\mathrm{T}}_{n}$
The proof strategy for Theorem 1.3 relies on establishing a connection between bounds on $\operatorname{Var}(e^{\lambda{\mathrm{T}}_{n}/2})$ and exponential concentration (see [4, Lemma 4.1]). To achieve the required variance bound, we employ the Falik-Samorodnitsky inequality (Lemma 4.1) to a martingale decomposition of the random variable $e^{\lambda{\mathrm{F}}_{m}}$, where ${\mathrm{F}}_{m}$ represents an averaged version of the passage time. This approach was first introduced by Benaïm and Rossignol in [4] and subsequently utilized by Damron, Hanson, and Sosoe in [16]. Finally, we estimate the tails of the true passage time ${\mathrm{T}}_{n}$ based on those of ${\mathrm{F}}_{m}$.
@@ 4.1. Variance bound via entropy inequality
$\forall i\geq 1,\quad\Delta_{i}:={\mathbb{E}}[G\mid{\mathcal{F}}_{i}]-{\mathbb{E}}[G\mid{\mathcal{F}}_{i-1}]={\mathbb{E}}[G(t_{e_{i}},t_{e_{i}^{c}})-G(t^{\prime}_{e_{i}},t_{e_{i}^{c}})\mid{\mathcal{F}}_{i}],$
where $t^{\prime}_{e_{i}}$ is an independent copy of $t_{e_{i}}$. We will bound the variance of $G$ based on the following entropy inequality by Falik and Samorodnitsky [22, Lemma 2.3].
@@
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]\geq\operatorname{Var}[G]\log\dfrac{\operatorname{Var}[G]}{\sum_{i=1}^{\infty}({\mathbb{E}}[|\Delta_{i}|])^{2}},$ (42)
@@
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]\leq C\sum_{i=1}^{\infty}{\mathbb{E}}[(G(b,t_{e_{i}^{c}})-G(a,t_{e_{i}^{c}}))^{2}].$ (43)
The detailed proof of this lemma can be found in [4, 17], or [18, Lemma 2.8].
4.2. Proof of Theorem 1.3 assuming Theorem 4.3
Instead of directly showing the subdiffusive concentration of ${\mathrm{T}}_{n}$}, we will employ a strategy inspired by Benjamini, Kalai, and Schramm in [5], known as the BKS trick. This approach involves proving the subdiffusive concentration for a geometric average of passage time, a concept used in [5, 2, 16, 17, 6]. Precisely, we define a spatial average of the first passage time,
${{\mathrm{F}}_{m}}:=\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}{\mathrm{T}}^{z},$ (44)
@@
$\forall z\in{\mathbb{Z}}^{d},\quad{\mathrm{T}}^{z}:={\mathrm{T}}(z,z+n{\mathbf{e}}_{1}),\qquad m:=\lfloor n^{1/4}\rfloor.$
@@
$\forall\,|\lambda|<\dfrac{1}{\sqrt{K}},\quad\operatorname{Var}\left[e^{\lambda{{\mathrm{F}}_{m}}}\right]\leq K\lambda^{2}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]<\infty,$ (45)
@@
The following result is a direct consequence of Theorem 4.3. We refer the reader to [4, Lemma 4.1] for a proof.
@@
$\mathbb{P}\left(\left|{{\mathrm{F}}_{m}}-{\mathbb{E}}[{{\mathrm{F}}_{m}}]\right|\geq\sqrt{\frac{n}{\log n}}\kappa\right)\leq c^{-1}\exp(-c\kappa)\quad\forall\kappa\geq 0.$ (46)
@@
where we recall that $z^{*}$ is the closest point of $z$ in the infinite cluster $\mathcal{C}$$_{\infty}$. The last two terms are bounded by $c_{1}^{-1}\exp\left(\tfrac{-c_{1}t}{\log^{2}n}\right)$, for some positive constants $c_{1}$ using Lemma 2.10, whereas by Lemma 23, the first term is bounded by $c_{2}^{-1}\exp(-c_{2}t)$ when $t\geq C\|x-y\|_{\infty}$ with some $C,c_{2}>0$. Hence, the result follows. ∎
@@
$\leq|{{\mathrm{F}}_{m}}-{\mathbb{E}}[{{\mathrm{F}}_{m}}]|+|{{\mathrm{T}}_{n}}-{{\mathrm{F}}_{m}}|.$ (48)
@@
$\mathbb{P}\left(|{{\mathrm{T}}_{n}}-{\mathbb{E}}[{{\mathrm{T}}_{n}}]|\geq 4M\right)\leq\mathbb{P}\left(|{{\mathrm{F}}_{m}}-{\mathbb{E}}[{{\mathrm{F}}_{m}}]|\geq 2M\right)+\mathbb{P}\left(|{{\mathrm{T}}_{n}}-{{\mathrm{F}}_{m}}|\geq 2M\right).$ (49)
@@
$|{{\mathrm{T}}_{n}}-{{\mathrm{F}}_{m}}|=\Big{|}{{\mathrm{T}}_{n}}-\dfrac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}{\mathrm{T}}^{z}\Big{|}$ $\leq\dfrac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}|{\mathrm{T}}(0,n{\mathbf{e}}_{1})-{\mathrm{T}}(z,z+n{\mathbf{e}}_{1})|$
$\leq\dfrac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}({\mathrm{T}}(0,z)+{\mathrm{T}}(n{\mathbf{e}}_{1},n{\mathbf{e}}_{1}+z)).$ (50)
@@
$\max_{z\in\Lambda_{m}}{\mathrm{T}}(0,z)\geq M\text{ or }\max_{z\in\Lambda_{m}}{\mathrm{T}}(n{\mathbf{e}}_{1},n{\mathbf{e}}_{1}+z)\geq M.$ (51)
@@
$\qquad\qquad\qquad=2\mathbb{P}\left(\max_{z\in\Lambda_{m}}{\mathrm{T}}(0,z)\geq M\right)\leq 2|\Lambda_{m}|\max_{z\in\Lambda_{m}}\mathbb{P}\left({\mathrm{T}}(0,z)\geq M\right),$ (52)
@@
$\max_{z\in\Lambda_{m}}\mathbb{P}\left({\mathrm{T}}(0,z)\geq M\right)\leq c^{-1}\exp(-cM/\log^{2}n),$ (53)
@@
Combining this with Corollary 4.4 and (49), it follows that
$\mathbb{P}\left(|{{\mathrm{T}}_{n}}-\mathbb{E}[{{\mathrm{T}}_{n}}]|\geq\sqrt{\frac{n}{\log n}}\kappa\right)\leq c^{-1}\exp(-c\kappa)\qquad\forall\kappa\geq 0,$ (54)
@@
Proof of Theorem 4.3. According to the Lemma 4.1, the variance bound is induced by two crucial factors: the estimate of total influence and total entropy. These keys are presented in the following results. From now on, we set
$G=e^{\lambda F_{m}},\quad\Delta_{i}={\mathbb{E}}[G|{\mathcal{F}}_{i}]-{\mathbb{E}}[G|{\mathcal{F}}_{i-1}],\quad\forall i\geq 1.$
@@
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq C\lambda^{2}n{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\quad\forall\,|\lambda|\leq\tfrac{1}{\log^{2(d+11)}n}.$ (55)
@@
$\operatorname{Var}\big{[}e^{\lambda{\mathrm{F}}_{m}}\big{]}\leq C\left(\log\frac{\operatorname{Var}[e^{\lambda{\mathrm{F}}_{m}}]}{C\lambda^{2}n^{(9-d)/8}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]}\right)^{-1}\lambda^{2}n{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}.$ (56)
We can assume that there exists $\lambda$ with $|\lambda|\leq\tfrac{1}{\log^{2(d+11)}n}$ such that
$\operatorname{Var}[e^{\lambda{\mathrm{F}}_{m}}]\geq C\lambda^{2}n^{15/16}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right],$ (57)
since otherwise there is nothing left to prove. By (56) and (57), there exists a constant $c>0$ such that for any $|\lambda|\leq\tfrac{1}{\log^{2(d+11)}n}$,